# Compositional Difference-in-Differences for categorical outcomes

### v2

@@ A.4. Proof of Proposition 3
I aim to show that the parallel growths assumption in compositional form is equivalent to a parallel trends condition in log-odds. Specifically, I prove the equivalence of the following two statements:
$(i)\quad\pi_{1,1}^{N}\ominus\pi_{0,1}^{N}=\pi_{1,0}^{N}\ominus\pi_{0,0}^{N}$
$(ii)\quad\textit{ for all }k=1,\dots,p-1,\log\left(\frac{\pi_{1,1}^{N}(c_{k})}{\pi_{1,1}^{N}(c_{p})}\right)-\log\left(\frac{\pi_{0,1}^{N}(c_{k})}{\pi_{0,1}^{N}(c_{p})}\right)=\log\left(\frac{\pi_{1,0}^{N}(c_{k})}{\pi_{1,0}^{N}(c_{p})}\right)-\log\left(\frac{\pi_{0,0}^{N}(c_{k})}{\pi_{0,0}^{N}(c_{p})}\right).$
@@
$\pi_{1}\ominus\pi_{2}:=\frac{1}{\sum_{k=1}^{p}\frac{\pi_{1}(c_{k})}{\pi_{2}(c_{k})}}\left(\frac{\pi_{1}(c_{1})}{\pi_{2}(c_{1})},\dots,\frac{\pi_{1}(c_{p})}{\pi_{2}(c_{p})}\right).$
I know that $\ell:\text{int}(\mathcal{S}^{p-1})\to\mathbb{R}^{p-1}$ defined by:
$\ell(\pi)=\left[\log\left(\frac{\pi(c_{1})}{\pi(c_{p})}\right),\dots,\log\left(\frac{\pi(c_{p-1})}{\pi(c_{p})}\right)\right].$
I already know that this map is a bijection. A key property of that transformation is that it linearizes the Aitchison geometry of the simplex. In particular, for any $\pi_{1},\pi_{2}\in\text{int}(\mathcal{S}^{p-1})$,
$\ell(\pi_{1}\ominus\pi_{2})=\ell(\pi_{1})-\ell(\pi_{2}).$
By definition of $\ominus$, the $k$-th component of $\pi_{1}\ominus\pi_{2}$ is proportional to $\pi_{1}(c_{k})/\pi_{2}(c_{k})$.
$\ell_{k}(\pi_{1}\ominus\pi_{2})$ $=\log\left(\frac{(\pi_{1}\ominus\pi_{2})(c_{k})}{(\pi_{1}\ominus\pi_{2})(c_{p})}\right)$
$=\log\left(\frac{\pi_{1}(c_{k})/\pi_{2}(c_{k})}{\pi_{1}(c_{p})/\pi_{2}(c_{p})}\right)$
$=\log\left(\frac{\pi_{1}(c_{k})}{\pi_{1}(c_{p})}\right)-\log\left(\frac{\pi_{2}(c_{k})}{\pi_{2}(c_{p})}\right)$
@@
$\pi_{1,1}^{N}\ominus\pi_{0,1}^{N}=\pi_{1,0}^{N}\ominus\pi_{0,0}^{N}\quad\Longleftrightarrow\quad\ell(\pi_{1,1}^{N}\ominus\pi_{0,1}^{N})=\ell(\pi_{1,0}^{N}\ominus\pi_{0,0}^{N}).$
@@ A.5. Proof of Theorem 2
Define the log cross-group ratio at time $t$ for category $c_{k}$:
$\Delta_{t}^{k}:=\log\left(\frac{q_{1,t}^{N}(c_{k})}{q_{0,t}^{N}(c_{k})}\right)=\log(q_{1,t}^{N}(c_{k}))-\log(q_{0,t}^{N}(c_{k})).$
The unobserved quantity of interest is $q_{1,1}^{N}(c_{k})$, or equivalently $\Delta_{1}^{k}$, since:
$q_{1,1}^{N}(c_{k})=\exp\left(\Delta_{1}^{k}+\log(q_{0,1}^{N}(c_{k}))\right).$
Thus, bounding $\Delta_{1}^{k}$ yields bounds on $q_{1,1}^{N}(c_{k})$. Assumption 3 states:
$\Delta_{1}^{k}\in\operatorname{Conv}\left\{\min_{t<0}\omega_{t}\Delta_{t}^{k},\;\max_{t<0}\omega_{t}\Delta_{t}^{k}\right\}=\left[\underline{\Delta}_{\omega}^{k},\;\overline{\Delta}_{\omega}^{k}\right],$
where
$\underline{\Delta}_{\omega}^{k}:=\min_{t<0}\omega_{t}\Delta_{t}^{k},\quad\overline{\Delta}_{\omega}^{k}:=\max_{t<0}\omega_{t}\Delta_{t}^{k}.$
Hence:
$b_{\min}^{k}=\exp\left(\underline{\Delta}_{\omega}^{k}+\log(q_{0,1}^{N}(c_{k}))\right)=\exp\left(\min_{t<0}\omega_{t}\log\left(\frac{q_{1,t}^{N}(c_{k})}{q_{0,t}^{N}(c_{k})}\right)+\log(q_{0,1}^{N}(c_{k}))\right),$
$b_{\max}^{k}=\exp\left(\overline{\Delta}_{\omega}^{k}+\log(q_{0,1}^{N}(c_{k}))\right)=\exp\left(\max_{t<0}\omega_{t}\log\left(\frac{q_{1,t}^{N}(c_{k})}{q_{0,t}^{N}(c_{k})}\right)+\log(q_{0,1}^{N}(c_{k}))\right).$
Let $q_{1,1}^{I}(c_{k})$ denote the observed outcome for group 1 in period 1 (i.e., the treated outcome). Since $q_{1,1}^{N}(c_{k})\in[b_{\min}^{k},b_{\max}^{k}]$, and the function $x\mapsto 1/x$ is decreasing for $x>0$, we have:
$\frac{q_{1,1}^{I}(c_{k})}{b_{\max}^{k}}-1\leq\text{GTT}(c_{k})\leq\frac{q_{1,1}^{I}(c_{k})}{b_{\min}^{k}}-1.$
Define total observed outcome: $S_{1,1}^{I}=\sum_{k=1}^{p}q_{1,1}^{I}(c_{k})$. The counterfactual total is $S_{1,1}^{N}=\sum_{k=1}^{p}q_{1,1}^{N}(c_{k})\in\left[\sum_{k}b_{\min}^{k},\sum_{k}b_{\max}^{k}\right]$. Then the aggregate treatment effect (e.g., ratio of totals) satisfies:
$\frac{S_{1,1}^{I}}{\sum_{k}b_{\max}^{k}}-1\leq\text{GTT}\leq\frac{S_{1,1}^{I}}{\sum_{k}b_{\min}^{k}}-1.$
Let $\pi_{1,1}^{I}=q_{1,1}^{I}/S_{1,1}^{I}$ be the observed share vector (in the simplex $\mathcal{S}^{p-1}$). The counterfactual share is $\pi_{1,1}^{N}=q_{1,1}^{N}/S_{1,1}^{N}$. The set of possible $\pi_{1,1}^{N}$ is:
$\left\{\pi\in\mathcal{S}^{p-1}:\pi_{k}=\frac{q_{k}}{\sum_{j}q_{j}},\;q_{k}\in[b_{\min}^{k},b_{\max}^{k}]\text{ for all }k\right\}.$
$\text{CTT}\in\left\{\pi\in\mathcal{S}^{p-1}:\pi=\pi_{1,1}^{I}\ominus s,\;s\in\ell^{-1}\left(\{r\in\mathbb{R}^{p-1}:\log(b_{\min}^{k})\leq r_{k}\leq\log(b_{\max}^{k})\}\right)\right\},$
A.6. Proof of Theorem 3
Strategy 1: Use the never-treated group ($G_{\infty}$) as control , it requires assumption 5: parallel growth (in logs) between each cohort $G_{g}$ and the never-treated group $G_{\infty}$.
Strategy 2: Use not-yet-treated groups as controls, it requires assumption 6: parallel growth (in logs) between cohort $G_{g}$ and any cohort $G_{s}$ that is not yet treated at time $t$ (i.e., $s>t$).
Because $\mathcal{G}$ excludes $\bar{g}$, for every $g\in\mathcal{G}$ and $t\geq g$, there exists at least one $s>t$ such that $G_{s}$ is not yet treated at time $t$ (since $s\leq\bar{g}$ and $\bar{g}>g$). Assumption 5 states: for all $g$ and $t\geq g$,
$\log q_{G_{g},t}^{N}-\log q_{G_{\infty},t}^{N}=\log q_{G_{g},t-1}^{N}-\log q_{G_{\infty},t-1}^{N}.$
This implies the difference is constant over time for $t\geq g$. In particular, it equals its value at t=g1t=g-1 (the last pre-treatment period):
$\log q_{G_{g},t}^{N}-\log q_{G_{\infty},t}^{N}=\log q_{G_{g},g-1}^{N}-\log q_{G_{\infty},g-1}^{N}.$
Rearranging:
$\log q_{G_{g},t}^{N}=\log q_{G_{g},g-1}^{N}+\log q_{G_{\infty},t}^{N}-\log q_{G_{\infty},g-1}^{N}.$
Exponentiating both sides gives the result. Assumption 6 states: for all $g\in\mathcal{G}$, and for all $s,t$ such that $t\geq g$ and $t<s\leq\bar{g}$,
$\log q_{G_{g},t}^{N}-\log q_{G_{s},t}^{N}=\log q_{G_{g},t-1}^{N}-\log q_{G_{s},t-1}^{N}.$
Again, this implies the log-difference is time-invariant from period g1g-1 onward (since $G_{s}$ is untreated at all times $\leq t<s$, so $q_{G_{s},\tau}^{N}=q_{G_{s},\tau}^{I}$ is observed for $\tau\leq t$). Thus, for any $t\geq g$ and any valid $s>t$,
$\log q_{G_{g},t}^{N}-\log q_{G_{s},t}^{N}=\log q_{G_{g},g-1}^{N}-\log q_{G_{s},g-1}^{N}.$
Rearranging and exponentiating gives the result.
@@ Appendix B Adding covariates
This section extends the main identification results to incorporate covariates. Let $X\in\mathcal{X}$ denote a vector of covariates, which may be endogenous. For this analysis, I restrict my attention to the case where $X$ is discrete. In this setting, I can redefine all relevant quantities by stratifying on $X$. This yields a set of new estimands, each conditional on a specific covariate profile $x$. Provided that the untreated potential quantities $q_{g,t}^{N}(x)$ are strictly positive for all groups, time periods, and values $x\in\mathcal{X}$, I can formulate the following conditional assumption:
@@
$\log\cdot\left(q_{1,1,x}^{N}\right)-\log\cdot\left(q_{1,0,x}^{N}\right)=\log\cdot\left(q_{0,1,x}^{N}\right)-\log\cdot\left(q_{0,0,x}^{N}\right).$
@@
$q_{1,1,x}^{N}=\exp\cdot\left(\log\cdot\left(q_{1,0,x}^{N}\right)+\log\cdot\left(q_{0,1,x}^{N}\right)-\log\cdot\left(q_{0,0,x}^{N}\right)\right)$ (15)
$q_{1,1}^{N}=\sum_{x\in\mathcal{X}}\mathbb{P}(X=x)q_{1,1,x}^{N}$ (16)
And the corresponding counterfactual probability mass can be obtained by: for each category $c_{k},k=1,\dots,p$
$\pi_{1,1,x}^{N}(c_{k})=\frac{\exp\left(\log\left(q_{1,0,x}^{N}(c_{k})\right)+\log\left(q_{0,1,x}^{N}(c_{k})\right)-\log\left(q_{0,0,x}^{N}(c_{k})\right)\right)}{\sum_{g=1}^{p}\exp\left(\log\left(q_{1,0,x}^{N}(c_{j})\right)+\log\left(q_{0,1,x}^{N}(c_{j})\right)-\log\left(q_{0,0,x}^{N}(c_{j})\right)\right)}$ (17)
$\pi_{1,1}^{N}(c_{k})=\sum_{x\in\mathcal{X}}\mathbb{P}(X=x)\pi_{1,1,x}^{N}(c_{k})$ (18)
@@ Appendix C Implementation of CoDiD and Inference procedure
For each group $g$ and each time period $t$, we have a random sample of $q_{g,t}$ individuals. Define $\hat{q}_{g,t}(c_{k})$ as the number of individuals in group $g$ at time $t$ in category $c_{k}$. The estimated counterfactual is given by:
$\hat{q}_{1,1}^{N}=\exp\cdot\left(\log\cdot\left(\hat{q}_{1,0}^{N}\right)+\log\cdot\left(\hat{q}_{0,1}^{N}\right)-\log\cdot\left(\hat{q}_{0,0}^{N}\right)\right)$
The total and shares are obtained using the estimated counterfactual quantities. The estimated treatment effects are given by:
$\hat{GTT}(c_{k})=\frac{\hat{q}^{I}_{1,1}(c_{k})}{\hat{q}^{N}_{1,1}(c_{k})}-1$
$\hat{GTT}=\frac{\hat{S}^{I}_{1,1}}{\hat{S}^{N}_{1,1}}-1$
$\hat{CTT}=\hat{\pi}^{I}_{1,1}\ominus\hat{\pi}^{N}_{1,1}$
For inference, I use the parametric bootstrap to approximate the sampling distribution of the parameters. The parametric bootstrap is a resampling technique designed for situations where the data-generating process can be modeled parametrically. Instead of drawing resamples directly from the observed data, the procedure assumes a parametric model for the data and uses estimated parameters to generate new synthetic samples. For a detailed treatment of the method, see Horowitz (2019). In my case, the procedure proceeds as follows. For each bootstrap iteration $b=1,\dots,B$, where $B$ is the total number of iterations, we have:
  • Resample quantities for each group-time pair $(g,t)$ using a multinomial distribution:
    $q_{g,t}^{(b)}(c_{1},\dots,c_{p})\sim\text{Multinomial}\left(S_{gt},\hat{\pi}_{g,t}(c_{1}),\dots,\hat{\pi}_{g,t}(c_{p})\right).$
  • Compute counterfactual total quantities and category shares based on the bootstrap samples.
  • Calculate the bootstrap treatment effects: $\hat{GTT}^{b}$ and $\hat{CTT}^{b}$.
The $95\%$ confidence interval for each category $c_{k}$ is then obtained from the $2.5\%$ and $97.5\%$ quantiles of the bootstrap distribution in each case.
@@ Appendix D General random utility model without IIA
But now, assume that the error terms jointly follow a Generalized Extreme Value (GEV) distribution,
$F(\varepsilon_{1},\dots,\varepsilon_{p})=\exp\left[-G\left(e^{-\varepsilon_{1}},\dots,e^{-\varepsilon_{p}}\right)\right],$
@@
$U^{N}_{gt}(c_{k})=V^{N}_{gt}(c_{k})+\varepsilon_{gt}(c_{k})=\mu^{N}_{gt}(c_{k})+\log\left(G_{k}(\mu^{N}_{gt})\right)-\log\left(G(\mu^{N}_{gt})\right)+\log\left(S^{N}_{gt}\right)+\varepsilon_{gt}(c_{k})$
@@
Model $G(\cdot)$ $\pi^{N}(c_{k})$
Logit $\sum_{j=1}^{p}e^{V^{N}(c_{j})}$ $\frac{e^{V^{N}(c_{k})}}{\sum_{j}e^{V^{N}(c_{j})}}$
Nested Logit $\sum_{m=1}^{M}\Big(\sum_{j\in\mathcal{B}_{m}}e^{V^{N}(c_{j})/\lambda_{m}}\Big)^{\lambda_{m}}$ $\frac{e^{V^{N}(c_{k})/\lambda_{m}}\Big(\sum_{j\in\mathcal{B}_{m}}e^{V^{N}(c_{j})/\lambda_{m}}\Big)^{\lambda_{m}-1}}{\sum_{r}\Big(\sum_{j\in\mathcal{B}_{r}}e^{V^{N}(c_{j})/\lambda_{r}}\Big)^{\lambda_{r}}}$
PCL $\sum_{r<s}\big(e^{V^{N}(c_{r})/\lambda_{rs}}+e^{V^{N}(c_{s})/\lambda_{rs}}\big)^{\lambda_{rs}}$ $\frac{\sum_{j\neq k}e^{V^{N}(c_{k})/\lambda_{kj}}\big(e^{V^{N}(c_{k})/\lambda_{kj}}+e^{V^{N}(c_{j})/\lambda_{kj}}\big)^{\lambda_{kj}-1}}{\sum_{r<s}\big(e^{V^{N}(c_{r})/\lambda_{rs}}+e^{V^{N}(c_{s})/\lambda_{rs}}\big)^{\lambda_{rs}}}$
GNL $\sum_{m=1}^{M}\Big(\sum_{j\in\mathcal{B}_{m}}(\alpha_{jm}e^{V^{N}(c_{j})})^{1/\lambda_{m}}\Big)^{\lambda_{m}}$ $\frac{(\alpha_{km}e^{V^{N}(c_{k})})^{1/\lambda_{m}}\Big(\sum_{j\in\mathcal{B}_{m}}(\alpha_{jm}e^{V^{N}(c_{j})})^{1/\lambda_{m}}\Big)^{\lambda_{m}-1}}{\sum_{r}\Big(\sum_{j\in\mathcal{B}_{r}}(\alpha_{jr}e^{V^{N}(c_{j})})^{1/\lambda_{r}}\Big)^{\lambda_{r}}}$
Let $G_{k}(\mathbf{y})=\frac{\partial G(\mathbf{y})}{\partial y_{k}}$, with $\mathbf{y}=(y_{1},\dots,y_{p})$. This derivative is crucial in determining how likely an individual is to choose alternative $c_{k}$ under the GEV model. If we evaluate it at $\mathbf{y}=\mathbf{e}^{\boldsymbol{\mu}^{N}}=\big(e^{V^{N}(c_{1})},\dots,e^{V^{N}(c_{p})}\big)$, the exponentiated systematic utilities, the term $G_{k}(\mathbf{e}^{\boldsymbol{\mu}^{N}})$ tells us how the attractiveness of option $c_{k}$ depends on the attractiveness of all the other available options. It adjusts for the fact that some choices are more closely related than others. For example, if two alternatives are close substitutes, a high utility for one will reduce the probability of choosing the other. Conversely, if options are weakly related, changing one has little effect on the probability of picking the other. Now, consider the following assumption: from the RUM (McFadden (1977)), we have:
$\pi_{gt}(c_{k})=\frac{\exp\Big(V^{N}_{gt}(c_{k})\Big)G_{k}(\mathbf{V^{N}_{gt}})}{G(\mathbf{V^{N}_{gt}})}=\frac{\exp\Big(\mu^{N}_{gt}(c_{k})\Big)G_{k}(\mathbf{\mu^{N}_{gt}})}{G(\mathbf{\mu^{N}_{gt}})}$
$q_{gt}^{N}(c_{k})=\pi_{gt}(c_{k})\cdot S^{N}_{gt}=\frac{\exp\Big(\mu^{N}_{gt}(c_{k})\Big)G_{k}(\mathbf{\mu^{N}_{gt}})}{G(\mathbf{\mu^{N}_{gt}})}\cdot S^{N}_{gt}.$
@@
$\log q_{gt}^{N}(c_{k})$ $=$ $\mu_{gt}^{N}(c_{k})-\log\left(G(\mathbf{\mu^{N}_{gt}})\right)+\log(S_{gt}^{N})+\log(G_{k}(\mathbf{\mu^{N}_{gt}}))$
$\mathbb{E}[U_{g1}(c_{k})]-\mathbb{E}[U_{g0}(c_{k})]$ $=$ $\log q_{g1}^{N}(c_{k})-\log q_{g1}^{N}(c_{k})$
From there, we have that assumption 2 is equivalent to parallel trends in expected utilities. The result for utilities differences also follow.

### v3

@@ A.4. Proof of Proposition 3
I aim to show that the parallel growth assumption in compositional form is equivalent to a parallel trends condition in log-odds. Specifically, I prove the equivalence of the following two statements:
$(i)\quad\pi_{1,1}^{0}\ominus\pi_{0,1}^{0}=\pi_{1,0}^{0}\ominus\pi_{0,0}^{0},$
$(ii)\quad\ell(\pi^{0}_{1,1})-\ell(\pi^{0}_{0,1})=\ell(\pi^{0}_{1,0})-\ell(\pi^{0}_{0,0})$
@@
$\pi_{1}\ominus\pi_{2}:=\frac{1}{\sum_{k=1}^{p}\frac{\pi_{k,1}}{\pi_{k,2}}}\left(\frac{\pi_{1,1}}{\pi_{1,2}},\dots,\frac{\pi_{p,1}}{\pi_{p,2}}\right).$
Also, we recall that:
$\ell(\pi)=\Bigg[\log\left(\frac{\pi_{1}}{\pi_{p}}\right),\dots,\log\left(\frac{\pi_{p-1}}{\pi_{p}}\right)\Bigg].$
This map is a bijection. A key property of that transformation is that it linearizes the Aitchison geometry of the simplex. In particular, for any $\pi_{1},\pi_{2}\in\text{int}(\mathcal{S}^{p-1})$,
$\ell_{k}(\pi_{1}\ominus\pi_{2})$ $=\log\left(\frac{(\pi_{1}\ominus\pi_{2})_{k}}{(\pi_{1}\ominus\pi_{2})_{p}}\right)$
$=\log\left(\frac{\pi_{k,1}/\pi_{2,k}}{\pi_{p,1}/\pi_{p,2}}\right)$
$=\log\left(\frac{\pi_{k,1}}{\pi_{p,1}}\right)-\log\left(\frac{\pi_{k,2}}{\pi_{p,2}}\right)$
@@
$\pi_{1,1}^{0}\ominus\pi_{0,1}^{0}=\pi_{1,0}^{0}\ominus\pi_{0,0}^{0}\quad\Longleftrightarrow\quad\ell(\pi_{1,1}^{0}\ominus\pi_{0,1}^{0})=\ell(\pi_{1,0}^{0}\ominus\pi_{0,0}^{0}).$
@@ A.5. Proof of Theorem 2
We proceed in four steps.
Step 1: Apply Assumption 3.
By the Robust Parallel Growth assumption (Assumption 3), for each category $k=1,\dots,p$,
$d_{k,1}=\log\!\left(q^{0}_{k,1,1}\right)-\log\!\left(q^{0}_{k,0,1}\right)\in\big[\,d^{\min}_{k},\;d^{\max}_{k}\,\big].$
Step 2: Bound the counterfactual quantity $q^{0}_{k,1,1}$.
Exponentiating the inequality yields
$\exp\!\left(d^{\min}_{k}\right)\leq\frac{q^{0}_{k,1,1}}{q^{0}_{k,0,1}}\leq\exp\!\left(d^{\max}_{k}\right).$
Multiplying through by $q^{0}_{k,0,1}>0$ gives
$\exp\!\left(d^{\min}_{k}+\log q^{0}_{k,0,1}\right)\leq q^{0}_{k,1,1}\leq\exp\!\left(d^{\max}_{k}+\log q^{0}_{k,0,1}\right).$
Define
$b_{k}^{\min}=\exp\!\left(d^{\min}_{k}+\log q^{0}_{k,0,1}\right),\quad b_{k}^{\max}=\exp\!\left(d^{\max}_{k}+\log q^{0}_{k,0,1}\right).$
Then
$q^{0}_{k,1,1}\in\big[\,b_{k}^{\min},\;b_{k}^{\max}\,\big],$
which establishes the first claim.
Step 3: Bound the total mass $S^{0}_{1,1}$.
The total counterfactual mass is $S^{0}_{1,1}=\sum_{k=1}^{p}q^{0}_{k,1,1}$. Since each $q^{0}_{k,1,1}$ is bounded independently, the sharp bounds on the sum are obtained by summing the individual extrema:
$\sum_{k=1}^{p}b_{k}^{\min}\leq S^{0}_{1,1}\leq\sum_{k=1}^{p}b_{k}^{\max}.$
Step 4: Characterize the counterfactual distribution $\pi^{0}_{1,1}$.
The counterfactual probability vector is defined by
$\pi^{0}_{1,1,k}=\frac{q^{0}_{k,1,1}}{S^{0}_{1,1}},\quad k=1,\dots,p.$
Because each $q^{0}_{k,1,1}\in[b_{k}^{\min},b_{k}^{\max}]$ and $S^{0}_{1,1}=\sum_{j=1}^{p}q^{0}_{j,1,1}$, the set of feasible distributions is precisely
$\left\{\pi\in\mathcal{S}^{p-1}:\pi_{k}=\frac{q_{k}}{\sum_{j=1}^{p}q_{j}},\;q_{k}\in[b_{k}^{\min},b_{k}^{\max}]\text{ for all }k\right\},$
where $\mathcal{S}^{p-1}$ denotes the (p1)(p-1)-dimensional simplex.
@@ Appendix B Adding covariates
This section extends the main identification results to incorporate covariates. Let $X\in\mathcal{X}$ denote a vector of covariates, which may be endogenous. For this analysis, I restrict my attention to the case where $X$ is discrete. In this setting, I can redefine all relevant quantities by stratifying on $X$. This yields a set of new estimands, each conditional on a specific covariate profile $x$. Provided that the untreated potential quantities $q_{g,t}^{0}(x)$ are strictly positive for all groups, time periods, and values $x\in\mathcal{X}$, I can formulate the following conditional assumption:
@@
$\log\left(q_{1,1,x}^{0}\right)-\log\left(q_{1,0,x}^{0}\right)=\log\left(q_{0,1,x}^{0}\right)-\log\left(q_{0,0,x}^{0}\right).$
@@
$q_{1,1,x}^{0}=\exp\left(\log\left(q_{1,0,x}^{0}\right)+\log\left(q_{0,1,x}^{0}\right)-\log\left(q_{0,0,x}^{0}\right)\right)$ (16)
$q_{1,1}^{0}=\sum_{x\in\mathcal{X}}\mathbb{P}(X=x)q_{1,1,x}^{0}$ (17)
And the corresponding counterfactual probability mass can be obtained by: for each category.
$\pi_{1,1}^{0}(c_{k})=\sum_{x\in\mathcal{X}}\mathbb{P}(X=x)\pi_{1,1,x}^{0}(c_{k})$ (18)
@@ Appendix C Implementation of CoDiD and Inference procedure
For each group $g$ and time period $t$, we observe a random sample of $q_{g,t}$ individuals and define $\hat{q}_{g,t}(c_{k})$ as the number of individuals in category $c_{k}$. The CoDiD framework estimates the counterfactual quantities for the treated group using
$\hat{q}_{1,1}^{0}=\exp\!\Big(\log(\hat{q}_{1,0}^{0})+\log(\hat{q}_{0,1}^{0})-\log(\hat{q}_{0,0}^{0})\Big),$
from which the counterfactual totals, shares, and treatment effects are derived:
$\hat{GTT}(c_{k})=\frac{\hat{q}^{1}_{1,1}(c_{k})}{\hat{q}^{0}_{1,1}(c_{k})}-1,\quad\hat{GTT}=\frac{\hat{S}^{1}_{1,1}}{\hat{S}^{0}_{1,1}}-1,\quad\hat{ATT}=\hat{\pi}^{1}_{1,1}-\hat{\pi}^{0}_{1,1}.$
For inference, I consider a parametric bootstrap Horowitz (2019) to approximate the sampling distribution of the parameters. Rather than resampling the observed data directly, the bootstrap assumes a multinomial model for the category counts and generates synthetic samples based on the estimated probabilities. Specifically, for each iteration $b=1,\dots,B$, category counts are resampled as
$(q_{1,g,t}^{(b)},\cdots,q_{p,g,t}^{(b)})\sim\text{Multinomial}\big(S_{g,t},\hat{\pi}_{1,g,t},\dots,\hat{\pi}_{p,g,t}\big),$
from which counterfactual totals, shares, and treatment effects ($\hat{GTT}^{b}$ and $\hat{ATT}^{b}$) are recomputed. The $95\%$ confidence intervals for each category are then obtained from the $2.5\%$ and $97.5\%$ quantiles of the bootstrap distributions.
@@ Appendix D General random utility model without IIA
Now, assume that the error terms jointly follow a Generalized Extreme Value (GEV) distribution,
$F(\varepsilon^{0}_{1},\dots,\varepsilon^{0}_{p})=\exp\left[-G\left(e^{-\varepsilon^{0}_{1}},\dots,e^{-\varepsilon^{0}_{p}}\right)\right],$
@@
$U^{0}_{k,g,t}=V^{0}_{k,g,t}+\varepsilon^{0}_{gt}(c_{k})=\mu^{0}_{k,g,t}-\log\left(G(\mu^{0}_{gt})\right)+\log\left(S^{0}_{g,t}\right)+\varepsilon^{0}_{gt}(c_{k})$
@@
Model $G(\cdot)$ $\pi^{0}(c_{k})$
Logit $\sum_{j=1}^{p}e^{V^{0}(c_{j})}$ $\frac{e^{V^{0}(c_{k})}}{\sum_{j}e^{V^{0}(c_{j})}}$
Nested Logit $\sum_{m=1}^{M}\Big(\sum_{j\in\mathcal{B}_{m}}e^{V^{0}(c_{j})/\lambda_{m}}\Big)^{\lambda_{m}}$ $\frac{e^{V^{0}(c_{k})/\lambda_{m}}\Big(\sum_{j\in\mathcal{B}_{m}}e^{V^{0}(c_{j})/\lambda_{m}}\Big)^{\lambda_{m}-1}}{\sum_{r}\Big(\sum_{j\in\mathcal{B}_{r}}e^{V^{0}(c_{j})/\lambda_{r}}\Big)^{\lambda_{r}}}$
PCL $\sum_{r<s}\big(e^{V^{0}(c_{r})/\lambda_{rs}}+e^{V^{0}(c_{s})/\lambda_{rs}}\big)^{\lambda_{rs}}$ $\frac{\sum_{j\neq k}e^{V^{0}(c_{k})/\lambda_{kj}}\big(e^{V^{0}(c_{k})/\lambda_{kj}}+e^{V^{0}(c_{j})/\lambda_{kj}}\big)^{\lambda_{kj}-1}}{\sum_{r<s}\big(e^{V^{0}(c_{r})/\lambda_{rs}}+e^{V^{0}(c_{s})/\lambda_{rs}}\big)^{\lambda_{rs}}}$
GNL $\sum_{m=1}^{M}\Big(\sum_{j\in\mathcal{B}_{m}}(\alpha_{jm}e^{V^{0}(c_{j})})^{1/\lambda_{m}}\Big)^{\lambda_{m}}$ $\frac{(\alpha_{km}e^{V^{0}(c_{k})})^{1/\lambda_{m}}\Big(\sum_{j\in\mathcal{B}_{m}}(\alpha_{jm}e^{V^{0}(c_{j})})^{1/\lambda_{m}}\Big)^{\lambda_{m}-1}}{\sum_{r}\Big(\sum_{j\in\mathcal{B}_{r}}(\alpha_{jr}e^{V^{0}(c_{j})})^{1/\lambda_{r}}\Big)^{\lambda_{r}}}$
Let $G_{k}(\mathbf{y})=\frac{\partial G(\mathbf{y})}{\partial y_{k}}$, with $\mathbf{y}=(y_{1},\dots,y_{p})$. This derivative is crucial in determining how likely an individual is to choose alternative $c_{k}$ under the GEV model. If we evaluate it at $\mathbf{y}=\mathbf{e}^{\boldsymbol{\mu}^{0}}=\big(e^{V^{0}(c_{1})},\dots,e^{V^{0}(c_{p})}\big)$, the exponentiated systematic utilities, the term $G_{k}(\mathbf{e}^{\boldsymbol{\mu}^{0}})$ tells us how the attractiveness of option $c_{k}$ depends on the attractiveness of all the other available options. It adjusts for the fact that some choices are more closely related than others. For example, if two alternatives are close substitutes, a high utility for one will reduce the probability of choosing the other. Conversely, if options are weakly related, changing one has little effect on the probability of picking the other. Now, consider the following assumption: from the RUM (McFadden (1977)), we have:
$\pi_{k,g,t}=\frac{\exp\Big(V^{0}_{k,g,t}\Big)G_{k}(\mathbf{V^{0}_{gt}})}{G(\mathbf{V^{0}_{gt}})}=\frac{\exp\Big(\mu^{0}_{k,g,t}\Big)G_{k}(\mathbf{\mu^{0}_{gt}})}{G(\mathbf{\mu^{0}_{gt}})}$
$q_{k,g,t}^{0}=\pi_{k,g,t}\cdot S^{0}_{g,t}=\frac{\exp\Big(\mu^{0}_{k,g,t}\Big)G_{k}(\mathbf{\mu^{0}_{gt}})}{G(\mathbf{\mu^{0}_{gt}})}\cdot S^{0}_{g,t}.$
@@
$\log q_{k,g,t}^{0}$ $=$ $\mu_{gt}^{0}(c_{k})-\log\left(G(\mathbf{\mu^{0}_{gt}})\right)+\log(S_{g,t}^{0})+\log(G_{k}(\mathbf{\mu^{0}_{gt}}))$
Here, we need an extra assumption which is :
$\log\left(G_{k}(\mu^{0}_{11})\right)-\log\left(G_{k}(\mu^{0}_{10})\right)=\log\left(G_{k}(\mu^{0}_{01})\right)-\log\left(G_{k}(\mu^{0}_{00})\right)$
This assumption is trivially satisfied in the multinomial logit case. With this additional assumption, we obtain that that assumption 2 is equivalent to parallel trends in expected utilities. The result for utilities differences also follow.